wiki:Status

Status

2016 February 27

C2/C3 were returned to the OAN/SPM at the end of 2015 and installed on the telescope at the end of January 2016. The instrument team is investigating some minor problems with mutual focus and exposure synchronization, but has determined that C2/C3 are working well enough for science observations.

2015 April 19

We do not have an estimate of the date of the return of C2/C3 from GSFC.

2015 January 17

The filters have been changed.

2014 November 24

C4 has been characterised.

2014 October 21

C4 has been installed and is confocal with C0 and C1.

2014 October 2

During the summer of 2014 the cryostat frequently lost vacuum spontaneously. Each time this happened, we lost several nights of observation as we pumped and cooled again. We initially thought the problem might have been due to elevated ambient temperatures in the dome, but the problem got worse even as the temperatures dropped in September. Our attempts to diagnose the problem on site, with the close participation of the staff of the OAN/SPM, have not been successful. Therefore, we will be returning the cryostat to GSFC for diagnostics and corrective maintenance. Since we do not fully understand the cause of problem, we cannot predict with confidence when the cryostat will return, but we would not be surprised if it were not available again until 2015 January.

The removal of the cryostat means that RATIR will no longer operate with the C2 (ZY) and C3 (JH) channels. However, the C0 (SDSS ugr and other filters) and C1 (SDSS i) channels will not be affected.

The filters in the C0 filter wheel in addition to SDSS ugr are Bessell UBV, Hα 6573/15, Hα 6690/91, and red continuum 6459/102.

In place of the cryostat, we will install a temporary channel designated “C4”. This will consist of an air-cooled FLI ML3200ME CCD with a fixed SDSS z' filter. The field will be about 2.5 × 1.8 arcmin, which is significantly smaller than the 5.4 × 5.4 arcmin fields of C0 and C1. We aim to install align the center of the C4 field with the centres of the C0/C1 fields. We will distribute details of the configuration and performance of this channel as soon as possible.

2014 August 10

We have reinstalled RATIR on the telescope.

2014 August 6

We have removed RATIR from the telescope in order to recoat the primary and secondary mirrors.

2014 March 26

We changed the filters in the C0 wheel. We now have SDSS ugr, BessellUBV, Hα 6573/15, Hα 6690/91, and red continuum 6459/102.

2013 May 30

We changed the filters in the C0 wheel. We now have SDSS ugr, Bessell UBV, and Hα 6573/15.

2013 March 1

We changed the filters in the C0 wheel. We now have SDSS ugr, Bessell V, and Strömgren-Crawford uvbyNW.

2013 January 10

We began observations again after the winter vacation.

2012 December 26

We mourn the loss of Leonid Georgiev, our friend and colleague on the RATIR instrument team, who died today after a short illness.

2012 December 18

The observatory closed for the winter vacation.

2012 December 15

We were due to warm up RATIR on 2012 December 17. However, during a severe snow storm on the night of 2012 December 14/15, snow entered the dome. To facilitate clearing the snow safely, we warmed up RATIR early.

2012 December 1

We have just serviced the instrument, flipping and realigning the cryogenic dichroic D3. That has eliminated the strong ghosts seen in C2.

We also installed "doughnuts" on the north side of the telescope and under the northern polar bearing, to support the helium hoses and prevent excessive bending. With that, we can now point over most of the sky normally available to the telescope.

2012 November 16

We installed 4270/48, 5050/50, and 6630/40 filters to be used as narrow BVR filters on very bright stars. These filters complement the SDSS ugr, Bessell UBV, and Ha 6573/15 filters already installed. We removed the Gunn i filter.

2012 August 5

The C0 channel was reinstalled. The filters installed in the filter wheel are SDSS ugr, Bessell UBV, Gunn i (for testing ghosting), and the Ha 6573/15 filter.

2012 May 7

The shutter on the C0 channel failed. The detector was returned to the vendor, FLI, to be repaired.

2012 April 25

The two infrared channels of RATIR were installed.

2012 February 18

The two optical channels of RATIR were installed. We are currently working to commission them and robotic operations of the telescope under RTS2.

2012 February 15

The interim instrument was removed prior to the installation of the two optical channels of RATIR.

2012 January 10

The telescope is operating robotically with an interim instrument consisting of a single CCD and a filter wheel.

We are still accepting proposals for this mode; see the call for proposals. Note that the complement of filters is different to that specified in the call for proposals.

Last modified 17 months ago Last modified on Apr 12, 2016 1:21:40 PM